# Palm distributions of superposed point processes for statistical inference

### v1

@@ 3.1 Minimum contrast estimation
$K_{\Phi_{1}+\Phi_{2}}(r)=\frac{1}{\rho}\left[\left\{K_{\Phi_{1}}(r)\rho_{1}+\rho_{2}|B(o,r)|\right\}\frac{\rho_{1}}{\rho}+\left\{\rho_{1}|B(o,r)|+K_{\Phi_{2}}(r)\rho_{2}\right\}\frac{\rho_{2}}{\rho}\right],$ (2)
where $|B(o,r)|$ denotes the volume of the ball in $\mathbb{X}$. Plugging (2) into (1), we need to solve an optimization problem to estimate the parameters of $\Phi_{1}$ and $\Phi_{2}$. Below, we describe an application where $\Phi_{1}$ is a repulsive point pattern and $\Phi_{2}$ is a random background noise. Specifically, $\Phi_{1}$ is assumed to be a dpp; the class of dpps is briefly recalled in the next section.
3.2 Determinantal point processes
dpps are a class of repulsive point processes that enjoy a high degree of analytical tractability; they have been used to model inhibitory structure in spatial datasets and for diversity-promoting sampling in machine learning. See, e.g., Lavancier2015; KuleszaTaskar2012 and references therein.
@@ 3.2 Determinantal point processes
where $C(x_{h},x_{w})_{h,w=1,\ldots,k}$ is the $k\times k$ matrix with entries $C(x_{h},x_{w})$. In particular, we assume that $\xi$ follows a Gaussian dpp (Lavancier2015) with kernel $C(x,y)=\rho_{\xi}\exp\{-\|(x-y)/\alpha\|^{2}\}$ parametrized by $(\rho_{\xi},\alpha)$, with $\rho_{\xi}<(\pi\alpha^{2})^{-1}$ to ensure the process is well-defined. From Lavancier2015, Ripley’s $K$-function of $\xi$ is
$K_{\xi}(r)=\pi r^{2}-\frac{\pi\alpha^{2}}{2}\left(1-{\rm e}^{-2r^{2}/\alpha^{2}}\right).$ (3)
While dpps allow for likelihood-based inference as discussed in Lavancier2015, this is typically numerically cumbersome due to a Fourier series expansion and the determinant of large matrices involved in the likelihood. Hence, in the spatstat package (BaddeleyTurner2005), the default way of fitting a dpp is through mce based on the $K$-function.
3.3 Fitting a corrupted determinantal point process via minimum contrast
Assume now to observe a realization of $\xi$ corrupted by a background noise, independent of $\xi$. Let $\Phi_{2}$ be a homogeneous Poisson point process on $R$ with intensity $\omega$, which models the corrupting noise. This setting naturally fits within our proposed modeling framework. Plugging (3) into (2), and recalling that for the homogeneous Poisson process $\Phi_{2}$ we have $K_{\Phi_{2}}(r)=\pi r^{2}$, the $K$-function for $\Phi=\xi+\Phi_{2}$ equals
$K_{\xi+\Phi_{2}}(r)=\pi r^{2}-\frac{\rho_{\xi}^{2}}{(\rho_{\xi}+\omega)^{2}}\frac{\pi\alpha^{2}}{2}\left(1-{\rm e}^{-2r^{2}/\alpha^{2}}\right).$ (4)
Let $\bm{x}=(x_{1},\ldots,x_{k})\in R^{k}$ denote the observed point pattern. The goal is to estimate the parameters of $\Phi=\xi+\Phi_{2}$, namely the intensity $\rho_{\xi}$ of the signal process, the repulsion parameter $\alpha$, and the noise intensity $\omega$. As customary, we estimate the overall intensity $\rho=\rho_{\xi}+\omega$ of $\Phi$ by $\hat{\rho}=k/|R|$, so that we are left with estimating only $\rho_{\xi}$ and $\alpha$. Following the discussion above, we minimize the objective function in (1), where $\hat{s}$ is the edge-corrected nonparametric estimator discussed by ripley1976second (see also MoWaBook03), and $s$ is the $K$-function in (4). Moreover, assuming that we observe data on a rectangular region $R=[a,b]\times[c,d]$, $r_{u}$ is set to one quarter of the smallest side length of $R$, following diggle2013statistical. For numerical purposes, we approximate the integral via numerical quadrature using Simpson’s rule. The minimization is performed via the bfgs algorithm using the Julia programming language.

### v2

@@ Acknowledgments
$K_{\Phi_{1}+\Phi_{2}}(r)=\frac{1}{\rho}\left[\left\{K_{\Phi_{1}}(r)\rho_{1}+\rho_{2}|B(o,r)|\right\}\frac{\rho_{1}}{\rho}+\left\{\rho_{1}|B(o,r)|+K_{\Phi_{2}}(r)\rho_{2}\right\}\frac{\rho_{2}}{\rho}\right],$ (S18)
where $|B(o,r)|=\pi r^{2}$ in $\mathbb{R}^{2}$.
$G$-function. The nearest neighborhood function (or $G$-function) for a stationary point process $\Phi$ is the cumulative distribution function of the distance between a typical point $o\in\Phi$ to the nearest other point of $\Phi$ (Baddeley et al., 2015). More specifically:
$G_{\Phi}(r)=P(\Phi_{o}^{!}(B(o,r))>0),$
and thanks to the stationarity, the function does not depend on the choice of $o$. As for the superposed point process $\Phi=\Phi_{1}+\Phi_{2}$, Theorem 1 leads us to the following
$\begin{split}G_{\Phi}(r)&=1-P(\Phi_{o}^{!}(B(o,r))=0)\\ &=1-\left[\frac{\rho_{1}}{\rho_{1}+\rho_{2}}(1-G_{\Phi_{1}}(r))(1-F_{\Phi_{2}}(r))+\frac{\rho_{2}}{\rho_{1}+\rho_{2}}(1-F_{\Phi_{1}}(r))(1-F_{\Phi_{2}}(r))\right]\end{split}$
where $F_{\Phi_{i}}(r)=P(\Phi_{i}(B(o,r))>0)$ is the contact distribution function of $\Phi_{i}$. Since $\Phi_{2}$ is a Poisson process, we have that $(1-F_{\Phi_{2}}(r))=e^{-\pi r^{2}\rho_{2}}$, then:
$G_{\Phi}(r)=1-\left[\frac{\rho_{1}}{\rho_{1}+\rho_{2}}(1-G_{\Phi_{1}}(r))e^{-\pi r^{2}\rho_{2}}+\frac{\rho_{2}}{\rho_{1}+\rho_{2}}(1-F_{\Phi_{1}}(r))e^{-\pi r^{2}\rho_{2}}\right].$ (S19)
AA-function. We finally concentrate on the $A$-function, introduced by Chiu (2008). This is defined as
$A_{\Phi}(s,r)=E\left(s^{\Phi_{o}^{!}(B(o,r))}\right)$
where $r\geq 0$ and $|s|<1$. For fixed $r$, this quantity represents the probability generating function of $\Phi_{o}^{!}(B(o,r))$. If the point process $\Phi$ is stationary, then the $A$-function does not depend on the choice of the point $o$, and so we can consider the origin as a reference point. A plain application of Theorem 1 and the fact that $\Phi_{2}$ is a Poisson point process lead us to the following formula for the $A$-function:
$A_{\Phi}(s,r)=\frac{\rho_{1}}{\rho_{1}+\rho_{2}}A_{\Phi_{1}}(s,r)\mathcal{G}_{\Phi_{2}(B(o,r))}(s)+\frac{\rho_{2}}{\rho_{1}+\rho_{2}}\mathcal{G}_{\Phi_{1}(B(o,r))}(s)\mathcal{G}_{\Phi_{2}(B(o,r))}(s)$
where $\mathcal{G}_{X}(s)$ denotes the probability generating function of a random variable $X$ evaluated at $s$. In addition, by exploiting the expression of $\mathcal{G}_{\Phi_{2}(B(o,r))}(s)$ for a Poisson random variable, we can write
$A_{\Phi}(s,r)=\frac{\rho_{1}}{\rho_{1}+\rho_{2}}e^{\rho_{2}\pi r^{2}(s-1)}A_{\Phi_{1}}(s,r)+\frac{\rho_{2}}{\rho_{1}+\rho_{2}}e^{\rho_{2}\pi r^{2}(s-1)}\mathcal{G}_{\Phi_{1}(B(o,r))}(s).$ (S20)
S5 Minimum contract estimation for stationary point processes
In Section S5.1 we give the details for the evaluation of the $A$-function and the $K$-function used in the experiment of Section 3.2 to fit the corrupted Matérn cluster process via mce. In Section S5.2, we provide another application of mce to fit a corrupted point process $\Phi=\Phi_{1}+\Phi_{2}$, where $\Phi_{1}$ is a repulsive point pattern, i.e., a determinantal point process, and $\Phi_{2}$ is a random background noise. The class of determinantal point processes is briefly recalled in Section S5.2.
S5.1 Fitting a corrupted Matérn cluster process: details for Section 3.2
Here we provide the explicit expressions of the $A$-function and the $K$-function used to fit the corrupted Matérn cluster process via minimum contrast estimation in Section 3.2. Having set
$\kappa(\xi;c)=\frac{1}{\pi R^{2}}{1}_{B(c,R)}(\xi),$
with $R>0$, recall that a Matérn cluster process is defined as
$\Phi_{1}\,|\,\Phi_{p}\sim\textsc{pp}\left(\int_{\mathbb{R}^{2}}\mu\kappa(\xi;c)\Phi_{p}(\mathrm{d}c)\mathrm{d}\xi\right),\quad\Phi_{p}\sim\textsc{pp}(\lambda\mathrm{d}c)$
where $\textsc{pp}(\lambda\mathrm{d}c)$ denotes a homogeneous Poisson process with intensity $\lambda$ on $\mathbb{R}^{2}$, and $\lambda,\mu>0$. The Matérn cluster process is a particular instance of the shot noise Cox process (Møller, 2003), thus the Palm distribution at a single point $x\in\mathbb{R}^{2}$ can be found in (Møller, 2003). We have
$\Phi_{1x}^{!}\stackrel{{\scriptstyle d}}{{=}}\Phi_{1}+\Phi_{\zeta_{x}},$ (S21)
where:
  • (i)
    $\Phi_{1}$ is a Matérn cluster process, distributed as the original one;
  • (ii)
    $\Phi_{\zeta_{x}}$ is a single cluster of points, where
    $\Phi_{\zeta_{x}}\,|\,\zeta_{x}\sim\textsc{pp}\left(\mu\kappa(\xi;\zeta_{x})\mathrm{d}\xi\right),\quad\zeta_{x}\sim f_{\zeta_{x}}(c)\propto\kappa(x;c).$
Finally, $\Phi_{1}$ and $\Phi_{\zeta_{x}}$ on the right hand side of (S21) are independent. It can be shown that the Matérn cluster process $\Phi_{1}$ is stationary with intensity $\rho_{1}=\lambda\mu$.
Now consider the corrupted point process $\Phi=\Phi_{1}+\Phi_{2}$, where $\Phi_{1}$ is a Matérn cluster process and $\Phi_{2}$ is a stationary Poisson process with intensity $\rho_{2}$. We first focus on the evaluation of the $A$-function, whose expression equals (S20). Since the process is stationary, the $A$-function does not depend on the specific point $x$, so we can use the origin $o$ as a reference point. Thanks to the representation of the Palm distribution (S21), the $A$-function of $\Phi_{1}$ equals
$\begin{split}A_{\Phi_{1}}(s,r)&=E\left[s^{\Phi_{1o}^{!}(B(o,r))}\right]=E\left[s^{\Phi_{1}(B(o,r))}\right]E\left[s^{\Phi_{\zeta_{o}}(B(o,r))}\right]\\ &=\mathcal{G}_{\Phi_{1}(B(o,r))}(s)E\left[s^{\Phi_{\zeta_{o}}(B(o,r))}\right].\end{split}$ (S22)
The expected value in (S22), can be evaluated as follows:
$E\left[s^{\Phi_{\zeta_{o}}(B(o,r))}\right]$ $=EE\left[s^{\Phi_{\zeta_{o}}(B(o,r))}\,|\,\zeta_{o}\right]=E\left[e^{\mu|B(o,r)\cap B(\zeta_{o},R)|(s-1)/(\pi R^{2})}\right]$
where we have used the fact that the inner expectation is the probability generating function of a Poisson random variable. Now, we can integrate out $\zeta_{o}$, whose density is $\kappa(o;c)$:
$E\left[s^{\Phi_{\zeta_{o}}(B(o,r))}\right]$ $=\int_{B(o,R)}e^{\mu|B(o,r)\cap B(c,R)|(s-1)/(\pi R^{2})}\frac{1}{\pi R^{2}}\mathrm{d}c.$
Therefore (S22) boils down to
$A_{\Phi_{1}}(s,r)=\mathcal{G}_{\Phi_{1}(B(o,r))}(s)\int_{B(o,R)}e^{\mu|B(o,r)\cap B(c,R)|(s-1)/(\pi R^{2})}\frac{1}{\pi R^{2}}\mathrm{d}c.$
We can plug in the previous expression in (S20) to obtain the $A$-function of the superposed point process $\Phi$:
$\begin{split}A_{\Phi}(s,r)&=\left[\frac{\rho_{1}}{\rho_{1}+\rho_{2}}\int_{B(o,R)}e^{\mu|B(o,r)\cap B(c,R)|(s-1)/(\pi R^{2})}\frac{1}{\pi R^{2}}\mathrm{d}c\right.\\ &\qquad\qquad\left.+\frac{\rho_{2}}{\rho_{1}+\rho_{2}}\right]e^{\rho_{2}\pi r^{2}(s-1)}\mathcal{G}_{\Phi_{1}(B(o,r))}(s),\end{split}$ (S23)
where it remains to evaluate the probability generating function of the Matérn cluster process $\Phi_{1}$. In order to do this, observe that the following representation holds true
$\Phi_{1}\,|\,\Phi_{p}\stackrel{{\scriptstyle d}}{{=}}\sum_{i\geq 1}\Phi_{1i},\quad\Phi_{1i}\,|\,\Phi_{p}\stackrel{{\scriptstyle\mbox{\scriptsize{\rm ind}}}}{{\sim}}\textsc{pp}(\mu\kappa(\xi;c_{i})\mathrm{d}\xi)$
where $\Phi_{p}=\sum_{i\geq 1}\delta_{c_{i}}$ is a homogeneous Poisson process with intensity $\lambda$. Thus, the previous representation entails
$\mathcal{G}_{\Phi_{1}(B(o,r))}(s)$ $=EE\left(s^{\Phi_{1}(B(o,r))}\,|\,\Phi_{p}\right)=E\left[\prod_{i\geq 1}E\left(s^{\Phi_{1i}(B(o,r))}\,|\,\Phi_{p}\right)\right]$
$=E\left[\prod_{i\geq 1}\exp\left\{\frac{\mu}{\pi R^{2}}|B(o,r)\cap B(c_{i},R)|(s-1)\right\}\right]$
$=E\left[\exp\left\{\int_{\mathbb{R}^{2}}\frac{\mu}{\pi R^{2}}|B(o,r)\cap B(c,R)|(s-1)\Phi_{p}(\mathrm{d}c)\right\}\right]$
$=\exp\left\{-\int_{\mathbb{R}^{2}}\left[1-\exp\left\{\frac{\mu}{\pi R^{2}}|B(o,r)\cap B(c,R)|(s-1)\right\}\right]\lambda\mathrm{d}c\right\}.$
The integral over $\mathbb{R}^{2}$ can be evaluated by a change of variables (polar coordinates) to get
$\mathcal{G}_{\Phi_{1}(B(o,r))}(s)$ $=\exp\left\{-2\pi\int_{0}^{R+r}\left[1-\exp\left\{\frac{\mu}{\pi R^{2}}I_{R,r}(\varrho)(s-1)\right\}\right]\lambda\varrho\mathrm{d}\varrho\right\},$
having defined $I_{R,r}(\varrho):=|B(o,r)\cap B((\varrho,0),R)|$. As a consequence, (S23) becomes
$\begin{split}&A_{\Phi}(s,r)=\left[\frac{\rho_{1}}{\rho_{1}+\rho_{2}}\int_{0}^{R}\exp\left\{\frac{\mu}{\pi R^{2}}I_{R,r}(\varrho)(s-1)\right\}\frac{2\varrho}{R^{2}}\mathrm{d}\varrho+\frac{\rho_{2}}{\rho_{1}+\rho_{2}}\right]\\ &\qquad\times e^{\rho_{2}\pi r^{2}(s-1)}\exp\left\{-2\pi\int_{0}^{R+r}\left[1-\exp\left\{\frac{\mu}{\pi R^{2}}I_{R,r}(\varrho)(s-1)\right\}\right]\lambda\varrho\mathrm{d}\varrho\right\},\end{split}$ (S24)
where
$I_{R,r}(\varrho)=\left\{\begin{array}[]{cc}\pi\min\{r,R\}^{2}&\text{if }\varrho\leq|r-R|\\ r^{2}\arccos{\frac{\varrho^{2}+r^{2}-R^{2}}{2\varrho R}}+R^{2}\arccos{\frac{\varrho^{2}+R^{2}-r^{2}}{2\varrho R}}-\frac{1}{2}\sqrt{p(\varrho,r,R)}&\text{otherwise }\end{array}\right$.
having set $p(\varrho,r,R):=(-\varrho+r+R)(\varrho+r-R)(\varrho-r+R)(\varrho+R+r)$.
As for the $K$-function, we can exploit (S18) to get
$K_{\Phi_{1}+\Phi_{2}}(r)=\frac{1}{\rho}\left[\left\{K_{\Phi_{1}}(r)\rho_{1}+\rho_{2}\pi r^{2}\right\}\frac{\rho_{1}}{\rho}+\left\{\rho_{1}\pi r^{2}+\pi r^{2}\rho_{2}\right\}\frac{\rho_{2}}{\rho}\right],$ (S25)
where we used the fact that $K_{\Phi_{2}}(r)=\pi r^{2}$, since $\Phi_{2}$ is a Poisson process, and that $|B(o,r)|=\pi r^{2}$. In addition, $K_{\Phi_{1}}(r)$ can be evaluated by resorting to (S21) to obtain
$K_{\Phi_{1}}(r)=\pi r^{2}+\frac{2\pi}{\lambda\pi^{2}R^{4}}\int_{0}^{R}\varrho I_{R,r}(\varrho)\mathrm{d}\varrho$
with some simple calculations. Putting the previous expression in (S25), we obtain
$\begin{split}K_{\Phi_{1}+\Phi_{2}}(r)=2\frac{\rho_{1}\rho_{2}\pi r^{2}}{\rho^{2}}+\pi r^{2}\frac{\rho_{1}^{2}}{\rho^{2}}+\pi r^{2}\frac{\rho_{2}^{2}}{\rho^{2}}+\frac{\rho_{1}^{2}}{\rho^{2}}\cdot\frac{2\pi}{\lambda\pi^{2}R^{4}}\int_{0}^{R}\varrho I_{R,r}(\varrho)\mathrm{d}\varrho.\end{split}$
S5.2 Fitting a corrupted determinantal point process
Determinantal point processes (dpps) are a class of repulsive point processes, see Lavancier et al. (2015); Kulesza and Taskar (2012) and references therein for a wide range of statistical application of dpps.
@@
where $C(x_{h},x_{w})_{h,w=1,\ldots,k}$ is the $k\times k$ matrix with entries $C(x_{h},x_{w})$. In particular, we assume that $\xi$ follows a Gaussian dpp (Lavancier et al., 2015) with kernel $C(x,y)=\rho_{\xi}\exp\{-\|(x-y)/\alpha\|^{2}\}$ parametrized by $(\rho_{\xi},\alpha)$, with $\rho_{\xi}<(\pi\alpha^{2})^{-1}$ to ensure the process is well-defined. From Lavancier et al. (2015), Ripley’s $K$-function of $\xi$ is
$K_{\xi}(r)=\pi r^{2}-\frac{\pi\alpha^{2}}{2}\left(1-{\rm e}^{-2r^{2}/\alpha^{2}}\right).$ (S26)
While dpps allow for likelihood-based inference as discussed in Lavancier et al. (2015), this is typically numerically cumbersome due to a Fourier series expansion and the determinant of large matrices involved in the likelihood. Hence, in the spatstat package (Baddeley and Turner, 2005), the default way of fitting a dpp is through mce based on the $K$-function.
Assume now to observe a realization of $\xi$ corrupted by a background noise, independent of $\xi$. Let $\Phi_{2}$ be a homogeneous Poisson point process on $R$ with intensity $\omega$, which models the corrupting noise. This setting naturally fits within the modeling framework of Section 3.2. Plugging (S26) into (S18), and recalling that for the homogeneous Poisson process $\Phi_{2}$ we have $K_{\Phi_{2}}(r)=\pi r^{2}$, the $K$-function for $\Phi=\xi+\Phi_{2}$ equals
$K_{\xi+\Phi_{2}}(r)=\pi r^{2}-\frac{\rho_{\xi}^{2}}{(\rho_{\xi}+\omega)^{2}}\frac{\pi\alpha^{2}}{2}\left(1-{\rm e}^{-2r^{2}/\alpha^{2}}\right).$ (S27)
Let $\underline{x}=(x_{1},\ldots,x_{k})\in R^{k}$ denote the observed point pattern. The goal is to estimate the parameters of $\Phi=\xi+\Phi_{2}$, namely the intensity $\rho_{\xi}$ of the signal process, the repulsion parameter $\alpha$, and the noise intensity $\omega$. As customary, we estimate the overall intensity $\rho=\rho_{\xi}+\omega$ of $\Phi$ by $\hat{\rho}=k/|R|$, so that we are left with estimating only $\rho_{\xi}$ and $\alpha$. Following the discussion above, we apply mce based on the minimization of the discrepancy between the $K$ function in (S27) and the edge-corrected nonparametric estimator discussed by Ripley (1976) (see also Møller and Waagepetersen (2003)). For numerical purposes, we approximate the integral via numerical quadrature using Simpson’s rule. The minimization is performed via the bfgs algorithm using the Julia programming language.